Carrie works dues south of her apartment. Her friend Sarah works due east of the apartment. They leave for work at the same time. By the time Carrie is 5 miles form their apartment, the distance between them is 1 mile more than Sarah’s distance from the apartment. How far from the apartment is Sarah?
Answer by greenestamps(13200)



The figure that can be drawn to display the given information is a right triangle with one leg 5 and hypotenuse 1 more than the other leg.

Many students will immediately recognize a 5-12-13 right triangle, giving...

ANSWER: Sarah is 12 miles from their apartment.

Using formal algebra and the Pythagorean Theorem...

5^2+x^2 = (x+1)^2
x^2+25 = x^2+2x+1
25 = 2x+1
2x = 24
x = 12

ANSWER: 12